# InfoMatch: Entropy Neural Estimation for Semi-Supervised Image Classification

### v1

@@ Abstract
Semi-supervised image classification, leveraging pseudo supervision and consistency regularization, has demonstrated remarkable success. However, the ongoing challenge lies in fully exploiting the potential of unlabeled data. To address this, we employ information entropy neural estimation to harness the potential of unlabeled samples. Inspired by contrastive learning, the entropy is estimated by maximizing a lower bound on mutual information across different augmented views. Moreover, we theoretically analyze that the information entropy of the posterior of an image classifier is approximated by maximizing the likelihood function of the softmax predictions. Guided by these insights, we optimize our model from both perspectives to ensure that the predicted probability distribution closely aligns with the ground-truth distribution. Given the theoretical connection to information entropy, we name our method InfoMatch. Through extensive experiments, we show its superior performance.
@@ 1 Introduction
Most existing SSL methods utilize two key strategies: pseudo supervision and consistency regularization. FixMatch Sohn et al. (2020) effectively combines these strategies. It utilizes weak augmented view to generate pseudo labels that serve as guidance for the prediction of strong augmented view. FlexMatch Zhang et al. (2021), FreeMatch Wang et al. (2023), and SoftMatch Chen et al. (2023) primarily focus on improving the quality and quantity of pseudo labels. These methods center on pseudo-label selection and leverage confidence metrics to assess uncertainty, labeling samples only when confidence surpasses a predefined threshold. Additionally, variants such as MixMatch Berthelot et al. (2019b) and ReMixMatch Berthelot et al. (2019a) delve into generating numerous novel samples through diverse data augmentation techniques, intending to introduce noise and transformations for enhanced model robustness.
@@
Our motivation does not focus on pseudo-label selection strategies and data augmentation techniques. Instead, we aim to efficiently exploit the potential of unlabeled data by adapting the entropy neural estimation from unsupervised representation learning Hjelm et al. (2019); Tian et al. (2023). The fundamental principle is to maximize the lower bound of mutual information between two augmented views Nowozin et al. (2016); Belghazi et al. (2018); Hjelm et al. (2019); Tian et al. (2023). These infomax approaches leverage all available unlabeled data, allowing the model to delve deeper into the inherent structure and patterns of the data, ultimately improving classification accuracy. We generate two strong augmentation views for all unlabeled data, utilizing the mutual information between these views as a lower bound on the dataset entropy. This strategy ensures the model consistently produces accurate output under varying input conditions, accurately capturing the most salient features of the data.
@@
Another motivation of this paper is to improve classifier accuracy by approximating the posterior probability. The probability of a sample belonging to a class is determined by the posterior of the classifier. A neural classifier models its last layer using a softmax function as a parameterized posterior, meaning the classifier’s predictive probability distribution approximates the ground-truth posterior distribution. We investigate the connection between posterior entropy and the likelihood function of its predictions. Our observation suggests that the upper bound of posterior entropy is effectively approximated by maximizing the softmax prediction’s likelihood. This finding is critical for advancing model optimization, ensuring the predicted probability distribution more accurately reflects the true distribution. We approximate the entropy of the ground-truth posterior. To apply the posterior entropy neural estimator to unlabeled samples, we employ pseudo supervision and weak-to-strong strategies in addition to the supervision loss of labeled data. We generate pseudo-labels through a weak augmentation view, guiding the prediction of two strong augmented views. Additionally, we introduce CutMix Yun et al. (2019) as a strong augmentation view to capture more non-deterministic target features.
@@
In our approach, we are motivated by two key objectives: estimating the entropy of the data and estimating the entropy of the ground-truth posterior. We propose a novel entropy-based methodology that integrates a data entropy neural estimator with a posterior entropy neural estimator. This combination prompts the model to thoroughly explore the intrinsic structure defined by entropy, achieved by maximizing the mutual information between augmented views of all unlabeled data. Moreover, we make the crucial observation that the upper bound on posterior entropy is effectively approximated by maximizing the predicted softmax likelihood function. This insightful observation is applied to unlabeled data in our approach through weak-to-strong, pseudo-supervision, and CutMix strategies. By approximating these entropies through gradient descent, our method progressively captures information about unlabeled data and model characteristics. Since our method primarily focuses on information entropy estimations, we aptly name it InfoMatch. The extensive experimental results consistently validate the effectiveness of our proposed approach, especially in scenarios where labeled data is scarce. Our dual-entropy-based method stands as a robust solution for fully exploiting the potential of unlabeled data, providing valuable insights into both the data structure and the posterior probabilities, thereby enhancing the performance of semi-supervised image classification.
@@
The main contributions are summarized as follows: 1) Our proposed InfoMatch effectively exploit the potential of unlabeled data. 2) Leveraging multiple objectives, InfoMatch approximates the data entropy and the ground-truth posterior entropy. By utilizing strong-to-strong constrastive, weak-to-strong pseudo supervision, and CutMix strategies, we have applied this approach to unlabeled data, enabling more efficient utilization of unlabeled data information. 3) Experiments across various semi-supervised learning benchmarks validate the superior performance of InfoMatch.
@@ 2 Related Works
Consistency regularization. Based on the smoothness assumption, consistency regularization suggests that minor perturbations applied to unlabeled samples should not result in substantial variations in their predictions. To ensure consistency, $\Pi$ Model Rasmus et al. (2015) introduces perturbations through data augmentation and dropout, aiming to maximize the similarity of predictions derived from two forward propagations of identical unlabeled sample. Temporal Ensembling Samuli and Timo (2017) employs a time series combination model to minimize the mean square error loss between current and historical predictions, thus streamlining forward reasoning. MeanTeacher Tarvainen and Valpola (2017) converts the exponential moving average of the prediction results into the model weight, penalizing the difference in predictions between student model and teacher model. In addition, Unsupervised Data Augmentation (UDA) Xie et al. (2020) extends the data augmentation method and employs specific target data augmentation algorithms for specific tasks. In contrast, Pseudo Label Lee (2013) works by generating artificial labels for unlabeled data and performing fully-supervised training using both labeled and pseudo-labeled data.
Holistic methods. Most SSL methods combine pseudo supervision with consistency regularization to achieve better performance. MixMatch Berthelot et al. (2019b) leverages unlabeled data by integrating both consistency regularization and entropy minimization. Building upon this, RemixMatch Berthelot et al. (2019a) introduces two innovative strategies, Distribution Alignment and Augmentation Anchor, to enhance its robustness and accuracy. FixMatch Sohn et al. (2020), which maintains the use of Augmentation Anchor, simplifies the number of strong augmemtations. And to improve the accuracy of pseudo labels, a fixed high threshold is established to eliminate unreliable labels. To alleviate class imbalance and address the issue of low utilization of unlabeled data in the early stage, FlexMatch Zhang et al. (2021) and FreeMatch Wang et al. (2023) take the model’s learning status and class learning difficulty into consideration to adaptively adjust the confidence threshold. SoftMatch Chen et al. (2023) uses a truncated Gaussian function to assign weights to samples based on their confidence, effectively addressing the inherent trade-off between the quantity and quality of pseudo labels.
@@
Mixing augmentation. The two prevalent mixing methods for images are Mixup Zhang et al. (2018) and CutMix Yun et al. (2019). Mixup generates novel training samples by combining two images and their corresponding labels linearly, aiming to facilitate the model in learning smoother decision boundaries. Conversely, CutMix swaps segments of one image with those of another, updating the corresponding labels simultaneously to encourage the model to comprehend and leverage local information within the image. Interpolation Consistency Training (ICT) Verma et al. (2022), MixMatch Berthelot et al. (2019b) and RemixMatch Berthelot et al. (2019a) combine Mixup with consistency regularization to expand the dataset by generating virtual “mixed” training samples, and utilize consistency regularization to ensure its generalization ability on mixed data. While FMixCutMatch Wei et al. (2021) combines Fourier space-based data cutting and data mixing to augment both labeled and unlabeled samples, thereby enhancing its performance and generalization. Furthermore, Ghorban et al. (2022) utilizes Mixup and CutMix simultaneously, aiming to blend the semantic information of three images and introduce enhanced perturbations into the data. To further enhance its attention to local information and improve its sensitivity to diverse position cues, we employ CutMix Yun et al. (2019) in InfoMatch.
In this paper, we introduce a novel approach that seamlessly integrates pseudo supervision, consistency regularization, and mixing strategies to improve its generalization capabilities. Additionally, we employ an entropy-based measure to assess its model evidence, encouraging it to delve deeper into the intrinsic structure of the dataset. This technical perspective enhances its comprehensive understanding of the underlying patterns within the dataset, ultimately leading to improved overall performance.
@@ 3 Entropy Neural Estimation
We employ an independent and identically distributed (i.i.d.) training dataset denoted as ${\mathcal{X}}={\mathcal{X}}_{l}\cup{\mathcal{X}}_{u}$, where ${\mathcal{X}}_{l}$ comprises labeled data pairs $({\bm{x}}_{i},{\bm{p}}_{i}),\forall\,i\in\{1,\ldots,n_{l}\}$, and ${\mathcal{X}}_{u}$ consists of unlabeled data ${\bm{x}}_{i},\forall\,i\in\{n_{l}+1,\ldots,n_{l}+n_{u}\}$. The sizes of the labeled and unlabeled datasets are denoted by $n_{l}$ and $n_{u}$, respectively, with ${n_{l}}\ll{n_{u}}$, and the total dataset size is $n=n_{l}+n_{u}$. Here, each ${\bm{x}}_{i}$ represents a data point, and ${\bm{p}}_{i}$ corresponds to its ground-truth class label, represented using a one-hot encoding. ${\mathcal{C}}=\{c_{1},\ldots,c_{k}\}$ denotes the class set. If ${\bm{x}}_{i}$ belongs to $c_{j}$ class, we have $p(c_{j}|{\bm{x}}_{i})=p_{ij}=1$
@@
According to the second term of the information bottleneck theory Tishby et al. (2000), $\min{\rm I}({\mathcal{X}};{\mathcal{Y}})-\beta{\rm I}({\mathcal{Y}};{\mathcal{C}})$, the neural network is mainly coding its softmax prediction $\sigma(c_{j}|{\bm{x}}_{i},\theta)$ to close to the ground-truth posterior $p(c_{j}|{\bm{x}}_{i})$
@@
Since ${\bm{x}}_{i}$ are sampled i.i.d., $p(c|{\bm{x}}_{i})$ corresponds point-to-point with ${\bm{x}}_{i}$ and is regarded as an i.i.d. set, the posterior distribution is $p({\mathcal{C}}|{\mathcal{X}})=\prod_{i=1}^{n}p(c|{\bm{x}}_{i})$ Then the expectation with respect to $p({\bm{x}})$ is approximated by a finite sum over ${\mathcal{X}}$, so that
@@ 3.2 Data Entropy Neural Estimation
${\rm I}({\mathcal{X}}^{(1)};{\mathcal{X}}^{(2)})={\rm div}_{\rm kl}\bigl{(}p({\bm{z}}^{(1)},{\bm{z}}^{(2)})\|p({\bm{z}}^{(1)})p({\bm{z}}^{(2)})\bigr{)}$ (6)
where ${\bm{z}}^{(1)}$ and ${\bm{z}}^{(2)}$ are the potential features corresponding to the two augmented views respectively.
@@ 3.3 Entropy Estimation for SSL
Given the modeling of the labeled data point $({\bm{x}},{\bm{p}})$ by the parameter $\theta$ to produce ${\bm{y}}=\sigma(c_{j}|{\bm{x}}_{i},\theta)$, we encode all labeled data ${\mathcal{X}}_{l}$, and the corresponding labeled data set is ${\mathcal{C}}_{l}$. During the training process, we strive to align the predictive coding of InfoMatch with the ground-truth coding, which essentially means modeling $y_{ij}$ to align closely with $p_{ij}$. In other words, we aim to approximate the entropy of the ground-truth posterior $-\ln p({\mathcal{C}}|{\mathcal{X}})$ with its negative log likelihood $-\ln\sigma({\mathcal{C}}|{\mathcal{X}},\theta)$. Referring to Lemma 1, the likelihood is
@@
where $p_{ij}$ is an element of $n_{l}\times k$ matrix of ground-truth matrix $P=[p_{ij}]$. The loss function is defined by taking the negative logarithm of the likelihood, resulting in the cross-entropy loss function:
@@
Referring to Lemma 2, we employ the strategy of contrastive learning by maximizing the lower bound of entropy for ${\mathcal{X}}_{u}$ We choose Jensen-Shannon divergence over Kullback-Leibler divergence, following the methodology in Nowozin et al. (2016). The estimation of the lower bound is
${\rm div}_{\rm js}\bigl{(}p({\bm{z}}^{(1)},{\bm{z}}^{(2)})\|p({\bm{z}}^{(1)})p({\bm{z}}^{(2)})\bigr{)}$
@@
where ${\rm div}_{\rm js}$ represents the Jensen-Shannon divergence.
Similar to Hjelm et al. (2019); Tian et al. (2023), we employ the contrastive loss, i.e., $\mathcal{L}_{\rm lower}$, for view-wise contrastive
@@
where $d(\cdot,\cdot)$ is the similarity score of pairwise logits, and $\mathcal{P}$ and $\mathcal{N}$ denote the sets of positive and negative pairs, respectively, i.e., ${\bm{z}}_{i}^{(1)}$ and ${\bm{z}}_{i}^{(2)}$ are belong to positive pairs $(i,i)\in\mathcal{P}$ while ${\bm{z}}_{i}^{(1)}$ and ${\bm{z}}_{j}^{(2)}$ are negative pairs $(i,j)\in\mathcal{N},\forall\,i\neq j$.

### v3

@@ Abstract
Semi-supervised image classification, leveraging pseudo supervision and consistency regularization, has demonstrated remarkable success. However, the ongoing challenge lies in fully exploiting the potential of unlabeled data. To address this, we employ information entropy neural estimation to utilize the potential of unlabeled samples. Inspired by contrastive learning, the entropy is estimated by maximizing a lower bound on mutual information across different augmented views. Moreover, we theoretically analyze that the information entropy of the posterior of an image classifier is approximated by maximizing the likelihood function of the softmax predictions. Guided by these insights, we optimize our model from both perspectives to ensure that the predicted probability distribution closely aligns with the ground-truth distribution. Given the theoretical connection to information entropy, we name our method InfoMatch. Through extensive experiments, we show its superior performance. The source code is available at https://github.com/kunzhan/InfoMatch.
@@ 1 Introduction
Most existing SSL methods utilize two key strategies: pseudo supervision and consistency regularization. FixMatch Sohn et al. (2020) effectively combines these strategies. It utilizes weak augmented view to generate pseudolabels that serve as guidance for the prediction of strong augmented view. FlexMatch Zhang et al. (2021), FreeMatch Wang et al. (2023), and SoftMatch Chen et al. (2023) primarily focus on improving the quality and quantity of pseudolabels. These methods center on pseudolabel selection and leverage confidence metrics to assess uncertainty, labeling samples only when confidence surpasses a predefined threshold. Additionally, variants such as MixMatch Berthelot et al. (2019b) and ReMixMatch Berthelot et al. (2019a) delve into generating numerous novel samples through diverse data augmentation techniques, intending to introduce noise and transformations for enhanced model robustness.
@@
Our motivation does not focus on pseudolabel selection strategies and data augmentation techniques. Instead, we aim to efficiently exploit the potential of unlabeled data by adapting the entropy neural estimation from unsupervised representation learning Hjelm et al. (2019); Tian et al. (2023). The fundamental principle is to maximize the lower bound of mutual information between two augmented views Nowozin et al. (2016); Belghazi et al. (2018); Hjelm et al. (2019); Tian et al. (2023). These infomax approaches leverage all available unlabeled data, allowing the model to delve deeper into the inherent structure and patterns of the data, ultimately improving classification accuracy. We generate two strong augmentation views for all unlabeled data, utilizing the mutual information between these views as a lower bound on the dataset entropy. This strategy ensures the model consistently produces accurate output under varying input conditions, accurately capturing the most salient features of the data.
@@
Another motivation of this paper is to improve classifier accuracy by approximating the posterior probability. The probability of a sample belonging to a class is determined by the posterior of the classifier. A neural classifier models its last layer using a softmax function as a parameterized posterior, meaning the classifier’s predictive probability distribution approximates the ground-truth posterior distribution. We investigate the connection between posterior entropy and the likelihood function of its predictions. Our observation suggests that the upper bound of posterior entropy is effectively approximated by maximizing the softmax prediction’s likelihood. This finding is critical for advancing model optimization, ensuring the predicted-probability distribution more accurately reflects the true distribution. We approximate the entropy of the ground-truth posterior. To apply the posterior entropy neural estimator to unlabeled samples, we employ pseudo supervision and weak-to-strong strategies in addition to the supervision loss of labeled data. We generate pseudolabels through a weak augmentation view, guiding the prediction of two strong augmented views. Additionally, we introduce CutMix Yun et al. (2019) as a strong augmentation view to capture more non-deterministic target features.
@@
In InfoMatch, we are motivated by two key objectives: estimating the entropy of the data and estimating the entropy of the ground-truth posterior. We propose a novel entropy-based methodology that integrates a data entropy neural estimator with a posterior entropy neural estimator. This combination prompts the model to thoroughly explore the intrinsic structure defined by entropy, achieved by maximizing the mutual information between augmented views of all unlabeled data. Moreover, we make the crucial observation that the upper bound on posterior entropy is effectively approximated by maximizing the predicted softmax likelihood function. This insightful observation is applied to unlabeled data in our approach through weak-to-strong, pseudo-supervision, and CutMix strategies. By approximating these entropies through gradient descent, our method progressively captures information about unlabeled data and model characteristics. Since our method primarily focuses on information entropy estimations, we aptly name it InfoMatch. The extensive experimental results consistently validate the effectiveness of InfoMatch, especially in scenarios where labeled data is scarce. Our dual-entropy-based method stands as a robust solution for fully exploiting the potential of unlabeled data, providing valuable insights into both the data structure and the posterior probabilities, thereby enhancing the performance of semi-supervised image classification.
@@
The main contributions are summarized as follows: 1) Our proposed InfoMatch effectively exploit the potential of unlabeled data. 2) Leveraging multiple objectives, InfoMatch approximates the data entropy and the ground-truth posterior entropy. By utilizing strong-to-strong constrastive, weak-to-strong pseudo supervision, and CutMix strategies, we apply this approach to unlabeled data, enabling more efficient utilization of unlabeled data information. 3) Experiments across various semi-supervised learning benchmarks validate the superior performance of InfoMatch.
@@ 2 Related Works
Consistency regularization. Based on the smoothness assumption, consistency regularization suggests that minor perturbations applied to unlabeled samples does not result in substantial variations in their predictions. To ensure consistency, $\Pi$ Model Rasmus et al. (2015) introduces perturbations through data augmentation and dropout, aiming to maximize the similarity of predictions derived from two forward propagations of identical unlabeled sample. Temporal Ensembling Samuli and Timo (2017) employs a time series combination model to minimize the mean square error loss between current and historical predictions, thus streamlining forward reasoning. MeanTeacher Tarvainen and Valpola (2017) converts the exponential moving average of the prediction results into the model weight, penalizing the difference in predictions between student model and teacher model. In addition, Unsupervised Data Augmentation (UDA) Xie et al. (2020) extends the data augmentation method and employs specific target data augmentation algorithms for specific tasks. In contrast, Pseudo Label Lee (2013) works by generating artificial labels for unlabeled data and performing fully-supervised training using both labeled and pseudo-labeled data.
Holistic methods. Most SSL methods combine pseudo supervision with consistency regularization to improve performance. MixMatch Berthelot et al. (2019b) leverages unlabeled data by integrating both consistency regularization and entropy minimization. Based on MixMatch, RemixMatch Berthelot et al. (2019a) introduces two innovative strategies, distribution alignment and augmentation anchor, to enhance its robustness and accuracy. FixMatch Sohn et al. (2020) which maintains the use of Augmentation Anchor, simplifies the number of strong augmemtations. To improve the accuracy of pseudolabels, a fixed high threshold is established to eliminate unreliable labels. To alleviate class imbalance and address the issue of low utilization of unlabeled data in the early stage, FlexMatch Zhang et al. (2021) and FreeMatch Wang et al. (2023) take the model’s learning status and class learning difficulty into consideration to adaptively adjust the confidence threshold. SoftMatch Chen et al. (2023) uses a truncated Gaussian function to assign weights to samples based on their confidence, addressing the inherent trade-off between the quantity and quality of pseudolabels.
@@
Mixing augmentation. The two prevalent mixing methods for images are mixup Zhang et al. (2018) and CutMix Yun et al. (2019). Mixup generates novel training samples by combining two images and their corresponding labels linearly, aiming to facilitate the model in learning smoother decision boundaries. Conversely, CutMix swaps segments of one image with those of another, updating the corresponding labels simultaneously to encourage the model to comprehend and leverage local information within the image. Interpolation consistency training Verma et al. (2022), MixMatch Berthelot et al. (2019b) and RemixMatch Berthelot et al. (2019a) combine mixup with consistency regularization to expand the dataset by generating virtual “mixed” training samples, and utilize consistency regularization to ensure its generalization ability on mixed data. While FMixCutMatch Wei et al. (2021) combines Fourier space-based data cutting and data mixing to augment both labeled and unlabeled samples, thereby enhancing its performance and generalization. Furthermore, Ghorban et al. (2022) utilizes mixup and CutMix simultaneously, aiming to blend the semantic information of three images and introduce enhanced perturbations into the data. To further enhance its attention to local information and improve its sensitivity to diverse position cues, we employ CutMix Yun et al. (2019) in InfoMatch.
@@ 3 Entropy Neural Estimation
We employ an independent and identically distributed (i.i.d.) training dataset denoted as ${\mathcal{X}}={\mathcal{X}}_{l}\cup{\mathcal{X}}_{u}$, where ${\mathcal{X}}_{l}$ comprises labeled data pairs $({\bm{x}}_{i},{\bm{p}}_{i}),\forall\,i\in\{1,\ldots,n_{l}\}$, and ${\mathcal{X}}_{u}$ consists of unlabeled data ${\bm{x}}_{i},\forall\,i\in\{n_{l}+1,\ldots,n_{l}+n_{u}\}$. The sizes of labeled and unlabeled datasets are denoted by $n_{l}$ and $n_{u}$, respectively, with ${n_{l}}\ll{n_{u}}$, and the total dataset size is $n=n_{l}+n_{u}$. Here, each ${\bm{x}}_{i}$ represents a data point, and ${\bm{p}}_{i}$ corresponds to its ground-truth class label, represented using a one-hot encoding. ${\mathcal{C}}=\{c_{1},\ldots,c_{k}\}$ denotes the class set. If ${\bm{x}}_{i}$ belongs to $c_{j}$ class, we have $p(c_{j}|{\bm{x}}_{i})=p_{ij}=1$
@@
According to the second term of the information bottleneck theory Tishby et al. (1999), $\min{\rm I}({\mathcal{X}};{\mathcal{Y}})-\beta{\rm I}({\mathcal{Y}};{\mathcal{C}})$, the neural network codes $\sigma(c_{j}|{\bm{x}}_{i},\theta)$ to close to $p(c_{j}|{\bm{x}}_{i})$ Here, ${\rm I}(\cdot,\cdot)$ denotes the mutual information, and ${\rm H}(\cdot)$ is the entropy.
@@
Since ${\bm{x}}_{i}$ are sampled i.i.d., $p(c|{\bm{x}}_{i})$ corresponds point-to-point with ${\bm{x}}_{i}$ and is regarded as an i.i.d. set, the posterior is $p({\mathcal{C}}|{\mathcal{X}})=\prod_{i=1}^{n}p(c|{\bm{x}}_{i})$ Then the expectation with respect to $p({\bm{x}})$ is approximated by a finite sum over ${\mathcal{X}}$, so that
@@ 3.2 Data Entropy Neural Estimation
${\rm I}({\mathcal{X}}^{(1)};{\mathcal{X}}^{(2)})={\rm div}_{\rm kl}\bigl{(}p({\bm{x}}^{(1)},{\bm{x}}^{(2)})\|p({\bm{x}}^{(1)})p({\bm{x}}^{(2)})\bigr{)}$ (6)
where ${\bm{x}}^{(1)}$ and ${\bm{x}}^{(2)}$ are two augmented data points.
@@ 3.3 Entropy Estimation for SSL
Given the modeling of labeled data point $({\bm{x}},{\bm{p}})$ by the parameter $\theta$ to produce ${\bm{y}}=\sigma(c_{j}|{\bm{x}}_{i},\theta)$, we encode all labeled data ${\mathcal{X}}_{l}$, and the corresponding labeled data set is ${\mathcal{C}}_{l}$. During the training process, we strive to align the predictive coding of InfoMatch with the ground-truth coding, which essentially means modeling $y_{ij}$ to align closely with $p_{ij}$. In other words, we aim to approximate the entropy of the ground-truth posterior $-\ln p({\mathcal{C}}|{\mathcal{X}})$ with its negative log likelihood $-\ln\sigma({\mathcal{C}}|{\mathcal{X}},\theta)$. Referring to Lemma 1, the likelihood is
@@
where $p_{ij}$ is an element of $n_{l}\times k$ matrix of ground-truth matrix $P=[p_{ij}]$. The loss is defined by taking the negative logarithm of the likelihood, resulting in the cross-entropy loss:
@@
Referring to Lemma 2, we employ the strategy of contrastive learning by maximizing the lower bound of entropy for ${\mathcal{X}}_{u}$ We choose Jensen-Shannon divergence over Kullback-Leibler divergence, following Nowozin et al. (2016). The lower bound $\mathcal{L}_{\rm{lower}}$ is derived from
${\rm div}_{\rm js}\bigl{(}p({\bm{x}}^{(1)},{\bm{x}}^{(2)})\|p({\bm{x}}^{(1)})p({\bm{x}}^{(2)})\bigr{)}$
@@
where ${\rm div}_{\rm js}$ represents the Jensen-Shannon divergence and $d(\cdot,\cdot)$ is the similarity score of pairwise logits.
Similar to Hjelm et al. (2019); Tian et al. (2023), we employ the view-wise contrastive loss, i.e., $\mathcal{L}_{\rm lower}$,
@@
where $\mathcal{P}$ and $\mathcal{N}$ denote positive and negative sets, respectively, i.e., ${\bm{z}}_{i}^{(1)}$ and ${\bm{z}}_{i}^{(2)}$ are belong to positive pairs $(i,i)\in\mathcal{P}$ while ${\bm{z}}_{i}^{(1)}$ and ${\bm{z}}_{j}^{(2)}$ are negative pairs $(i,j)\in\mathcal{N},\forall\,i\neq j$.